NW Furniture is considering adding a cedar picnic table to the line of furniture. NW Furniture estimates that it will sell for $50 to distributors. They also estimate that the fixed costs of producing that table will be $13,000 and that the variable costs per table will be $22.

How many table will they need to sell to break even?

Find the total sales they need to break even?



1
Expert's answer
2021-09-27T06:02:30-0400

"Contribution = Sale - Variable \space cost\\

= \$ 50 - \$22\\

= \$ 28\\

P/V Ratio = \frac{Contribution}{sales }\times 100\\

= \frac{28 }{ 50} \times 100\\

= 56\%"

A) Number of table to be sold for breakeven = Fixed cost "\div" Contribution

"= \frac{13000 }{ 28}\\

= 464.29\\

= 464 \space tables\space ( Approx)"

B) Total sales they need to break-even = Fixed cost "\div" PV Ratio

"= \frac{13000 }{ 56\%}

= \$ 23214.29\\

= \$ 23214 \space (Approx)"